solving eq
y^2(x-y)p+x^2(y-x)q-z(x^2+y^2)=0
1
Expert's answer
2020-06-15T18:21:19-0400

#Q122028

Given PDE is

y2(x-y)p+x2(y-x)=(x2+y2)z

Now we can find out its general solution by using Lagrange's method

"\implies" dx/y2(x-y) = dy/x2(y-x) = dz/ (x2+y2)

For finding 1st pair

dx-dy/(x2+y2)(x-y) = dz/(x2+y2)

So, dx-dy/(x-y) = dz ...(1)

integrate equation 1st we get

ln(x-y)=z+lnc1 "\implies" ln(x-y/c1)=z "\implies" (x-y)/c1=ez "\implies" (x-y)/ez .....(*)

Now let us find another pair

dx+ydz/xy(x+y) = dy+xdz/xy(x+y)

"\implies" dx+ydz = dy+xdz ....(2)

Taking integration on both side of (2 ), we get

x+yz =y+xz+c2 "\implies" (x-y)-z(x-y)=c2 ......(**)

Therefore general solution is given by (*) and (**)

So, general solution is : "\Phi" [(x-y)(1-z)]=(x-y)/ez Ans.